# KnowGPT: Black-Box Knowledge Injection for Large Language Models

### 11 Dec 2023 v1

@@ Abstract
Generative Large Language Models (LLMs), such as ChatGPT, offer interactive APIs that can answer common questions at a human-expert level. However, these models often give inaccurate or incorrect responses when faced with questions requiring domain-specific or professional-specific knowledge not covered in their training corpus. Furthermore, many state-of-the-art LLMs are not open-source, making it challenging to inject knowledge with model APIs only. In this work, we introduce KnowGPT, a black-box knowledge injection framework for LLMs in question answering. KnowGPT leverages deep reinforcement learning (RL) to extract relevant knowledge from Knowledge Graphs (KGs) and use Multi-Armed Bandit (MAB) to construct the most suitable prompt for each question. Our extensive experiments on three benchmark datasets showcase that KnowGPT significantly enhances the existing methods. Notably, KnowGPT achieves an average improvement of 23.7% over ChatGPT and an average improvement of 2.9% over GPT-4. Additionally, KnowGPT attains a 91.6% accuracy on the OpenbookQA official leaderboard, which is comparable to human-level performance.
11footnotetext: Equal contribution.
@@ 1 Introduction
Generative large language models (LLMs) have surprised the world with their superior performance Kung et al. (2023); Zha et al. (2023), especially with the emergence of ChatGPT and GPT-4 OpenAI (2023). Nonetheless, LLMs are often criticized for their limited factual knowledge and propensity to produce hallucinations, wherein the model fabricates incorrect statements on tasks beyond their knowledge and perception Amaro et al. (2023); Shen et al. (2023); Gravel et al. (2023). Consider an ecological domain-specific question from OpenbookQA Mihaylov et al. (2018) in Figure 1. ChatGPT erroneously provides the response “energy” when asked about the portion of nutrients. This inaccuracy could stem from its potential lack of knowledge of carbs and their relationship to nutrients.
A promising avenue for addressing the above issue entails the integration of Knowledge Graphs (KGs) into LLMs. KGs, such as Yago Suchanek et al. (2007), Freebase Bollacker et al. (2008), and ConceptNet Speer et al. (2017) represent relationships among real-world entities in a structured form as triples (head, relation, tail). The enormous factual knowledge stored in KGs holds the potential to significantly enhance the accuracy of LLMs’ responses. For instance, in Figure 1, ChatGPT could correct itself by leveraging the related background knowledge in ConceptNet Speer et al. (2017).
Recently, various white-box knowledge injection methods have been proposed to incorporate KGs into LLMs. These methods assume that we know everything about LLMs, such as model architectures and parameters. Many algorithms have been proposed to integrate KGs into LLMs Pan et al. (2023). Early studies directly concatenate the entities from KGs with textual sentences as the input to train LLMs based on cross-modal representation alignment Sun et al. (2020); Liu et al. (2020b). Later research incorporates KGs at an implicit level by either combining text representation and KG embedding with the attention mechanism Feng et al. (2020); Yasunaga et al. (2021); Lin et al. (2019); Dong et al. (2023a) or designing various fusion methods to integrate KGs and texts through a tailored encoder Zhang et al. (2019); Su et al. (2021).
@@
Figure 1: A real-world question from OpenbookQA. ChatGPT could effectively correct the answer given the scientific reasoning background from ConceptNet (blue: question concepts, red: answers, grey: entities not present in questions).
@@
Unfortunately, many state-of-the-art LLMs are confined to a black-box role in practical applications. For instance, ChatGPT and GPT-4 exclusively grant access through their APIs, which means we can only retrieve model responses by submitting textual inputs, with model specifics inaccessible. This lack of access prevents us from employing the aforementioned white-box knowledge injection techniques. Even though white-box approaches could be applied to open-source LLMs, such as BLOOM Scao et al. (2022) and LLaMA Touvron et al. (2023), they often incur significant computation costs due to updating model weights Liu et al. (2022). Thus, we ask: Can we develop a black-box knowledge injection framework that can efficiently and effectively integrate KGs into LLMs with APIs only?
@@
Achieving this goal is nontrivial because of two challenges in constructing model inputs, or prompts. Identifying the most relevant knowledge is difficult. Real-world KGs often consist of millions of triples, whereas LLMs are typically restricted by limited input lengths (e.g., 2048 tokens for ChatGPT and 4096 tokens for GPT-4). Hence, careful selection of the most informative triples from KGs becomes essential. Effectively encoding KG knowledge is hard. It is observed that even minor variations in prompts conveying the same semantic meaning can yield drastically different responses from LLMs OpenAI (2023). As a result, a customized approach to encoding factual knowledge from extracted KGs for each question is often required to achieve the best performance.
@@
In this work, we propose KnowGPT, a black-box knowledge injection framework for LLMs in question answering. To address challenge , we leverage deep reinforcement learning (RL) to extract paths from source entities mentioned in the question to the target entities within the potential answers. To encourage the agent to discover more informative paths, we devise a tailored reward scheme that promotes the reachability, context-relatedness, and conciseness of the extracted paths. Then, a policy network is trained to maximize the reward using training questions and applied to unseen questions. To tackle challenge , we introduce a prompt construction strategy based on Multi-Armed Bandit (MAB). Given several path extraction strategies and prompt templates, a MAB is learned to select the most effective combination for each question by balancing exploration and exploitation. The learned MAB is then applied to new questions to select path extraction strategies and prompt templates automatically. Our main contributions are summarized as follows:
  • Formally define the problem of black-box knowledge injection for LLMs, which integrates KGs into LLMs with model APIs only.
  • Propose KnowGPT, a black-box knowledge injection framework for LLMs, which leverages deep RL for path extraction in KGs and MAB for prompt construction.
  • Implement KnowGPT upon two real-world KGs, ConceptNet Speer et al. (2017) and USMLE Yasunaga et al. (2021), with ChatGPT APIs. Experiments on three QA benchmarks show that KnowGPT outperforms the state-of-the-art baselines by a large margin. Notably, KnowGPT attains a 91.6% accuracy on the OpenbookQA leaderboard, which is comparable to human performance.
@@
Figure 2: The overall architecture of our proposed black-box knowledge injection framework, i.e., KnowGPT. Given the question context and answer choices, we retrieve a question-specific subgraph from the real-world KG. Path Extraction is first dedicated to searching for the most informative and concise reasoning background subject to the context. Then the prompt translation module is optimized to prioritize the combination of knowledge and formats subject to the given question.
@@ 2 Problem Statement
We formally define the problem of black-box knowledge injection for LLMs in complex question answering. We represent each question as a question context $\mathcal{Q}=\{\mathcal{Q}_{s},\mathcal{Q}_{t}\}$, where $\mathcal{Q}_{s}=\{e_{1},e_{2},...,e_{m}\}$ is a set of $m$ source entities, and $\mathcal{Q}_{t}=\{e_{1},e_{2},...,e_{n}\}$ is a set of $n$ target entities. Following prior work Feng et al. (2020); Yasunaga et al. (2022), $\mathcal{Q}_{s}$ is extracted by concept recognition, and we assume it is given in our problem. Similarly, each target entity in $\mathcal{Q}_{t}$ is extracted from a corresponding candidate answer. We denote an LLM as $f$, a real-world KG as $\mathcal{G}$, which consists of triples (head entity, relation, tail entity), denoted as $(h,r,t)$. In our setting, we only have access to the APIs of $f$. However, we can employ open-source lightweight language models (not $f$), like Bert-Base Devlin et al. (2018), to obtain text embeddings. Using the above notations, we describe our problem below.
Given a question context $\mathcal{Q}$, an LLM $f$, and a KG $\mathcal{G}$, we aim to learn a prompting function $f_{\text{prompt}}(\mathcal{Q},\mathcal{G})$, which generates a prompt x that incorporates the context of $\mathcal{Q}$ and the factual knowledge in $\mathcal{G}$, such that the prediction of the LLM $f(\textbf{x})$ can output the correct answers for $\mathcal{Q}$.
@@ 3 KnowGPT Framework
Learning the prompting function $f_{\text{prompt}}(\mathcal{Q},\mathcal{G})$ involves two challenges, i.e., what knowledge should be used in $\mathcal{G}$, and how to construct the prompt. To address these challenges, we present KnowGPT, which extracts subgraphs (paths) with deep RL and then constructs the prompt with MAB. An overview of our framework is shown in Figure 2.
3.1 Path Extraction with Deep Reinforcement Learning
Intuitively, the relevant reasoning background lies in a question-specific subgraph $\mathcal{G}_{\text{sub}}$ that contains all the source entities $\mathcal{Q}_{s}$, target entities $\mathcal{Q}_{t}$, and their neighbors. An ideal subgraph $\mathcal{G}_{\text{sub}}$ is expected to have the following properties: $(i)$ $\mathcal{G}_{\text{sub}}$ encompasses as many source and target entities as possible, $(ii)$ the entities and relations within $\mathcal{G}_{\text{sub}}$ exhibit a strong relevance to the question context, and $(iii)$ $\mathcal{G}_{\text{sub}}$ is concise with little redundant information such that it can be fed into LLMs with limited input lengths.
However, it is challenging to find such a $\mathcal{G}_{\text{sub}}$ since extracting a subgraph is NP-hard. To effectively and efficiently find a satisfactory $\mathcal{G}_{\text{sub}}$, we develop a tailored path extraction method, named $\mathcal{P}_{\text{RL}}$, that employs deep RL to sample reasoning paths in a trial-and-error fashion. Specifically, we assume $\mathcal{G}_{\text{sub}}$ is constructed based on a set of reasoning paths $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{m}\}$, where each path $\mathcal{P}_{i}=\{(e_{i},r_{1},t_{1}),(t_{1},r_{2},t_{2}),...,(t_{|\mathcal{P}_{i}|-1},r_{|\mathcal{P}_{i}|},t_{|\mathcal{P}_{i}|})\}$ is a path in $\mathcal{G}$ starting from the $i$-$th$ source entity in $\mathcal{Q}_{s}$, and $|\mathcal{P}_{i}|$ is the path length. $\mathcal{G}_{\text{sub}}$ encompasses all the entities and relations appeared in $\mathcal{P}$. We model the sampling of each reasoning path as a Markov Decision Process (MDP) with state, action, transition, and reward, defined as follows.
@@ 3.1 Path Extraction with Deep Reinforcement Learning
    where $\boldsymbol{e}_{t}$ and $\boldsymbol{e}_{target}$ are the embedding vectors of the current entity and the target entity. To get the initial node embeddings for entities extracted from the background KG, we adopt the approach proposed by the previous study Feng et al. (2020). Specifically, we transform knowledge triples from the KG into sentences and feed them into pre-trained LM to get node embeddings.
@@ 3.1.1 Context-relatedness Auxiliary Reward
where $\boldsymbol{c}$ is the embedding of context $\mathcal{Q}$ we obtained from a pre-trained LM Devlin et al. (2018) and the embedding of path $\mathcal{P}_{i}$ is the average of the embeddings of all the entities and relations we have walked through till $i$, i.e., $Avg(\boldsymbol{e}_{source}+\boldsymbol{re}_{1}...+\boldsymbol{e}_{i})$, where $i\leq length(\mathcal{P}_{target})$. This step-by-step reward scheme provides rewards before the target is reached.
@@ 3.1.3 Training Policy Network
3.2 Prompt Construction with Multi-armed Bandit
@@
Suppose we have several path extraction strategies $\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{m}\}$ and several candidate formats =$\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},...,\mathcal{F}_{n}\}$. Each path extraction strategy $\mathcal{P}_{i}$ is a method for selecting a subgraph given a question context, such as the RL-based strategy discussed above. Every prompt template $\mathcal{F}_{j}$ represents a mechanism to transform the triples within the subgraph into a prompt for an LLM prediction.
@@
To capture the context-aware correlation between questions and different combinations of knowledge and prompt formats, we formulate the selection mechanism of MAB with an expectation function $E(\cdot)$. It adaptively measures the potential expectation of a combination for different questions.
$\small E(\mathcal{Q}|\mathcal{P}\mathcal{F}_{(i)})=\boldsymbol{c}\times\bm{\alpha}_{(i)}+\beta_{(i)}.\vspace{1mm}$ (6)
@@
Empirically maximizing $\boldsymbol{c}\times\bm{\alpha}_{i}$ could encourage exploitation Chen et al. (2019); Dong et al. (2023b) for the best combination, we could effectively update $\bm{\alpha}_{(i)}$ via modeling the correlations between the context embedding of the anchor question ${\bf c}_{i}$ and all the previously selected contexts ${\bf C}_{(i)}$ for particular combination $\mathcal{P}\mathcal{F}_{(i)}$ in former $k$ steps, and the rewards $\textbf{r}_{pf}^{(i)}$ obtained from the selection of the current combination. Concretely, the $\bm{\beta}^{(b)}$ is updated as:

### 13 Mar 2024 v4

# KnowGPT: Knowledge Injection for Large Language Models

@@ Abstract
Generative Large Language Models (LLMs), such as ChatGPT and GPT-4, offer interactive APIs that can answer common questions at the human-expert level. However, these models often give inaccurate responses when faced with questions requiring domain-specific or professional-specific knowledge not covered in their training corpus. Furthermore, many state-of-the-art LLMs are not open-source, making it challenging to inject knowledge with model APIs only. In this work, we introduce KnowGPT, a black-box knowledge injection framework for LLMs in question answering. KnowGPT leverages deep reinforcement learning (RL) to extract relevant knowledge from Knowledge Graphs (KGs) and use Multi-Armed Bandit (MAB) to construct the most suitable prompt for each question. Our extensive experiments on three benchmark datasets showcase that KnowGPT significantly enhances the existing methods. Notably, KnowGPT achieves an average improvement of 23.7% over ChatGPT and an average improvement of 2.9% over GPT-4. Additionally, KnowGPT attains a 91.6% accuracy on the OpenbookQA official leaderboard, which is comparable to human-level performance. The code will be open-sourced.
KnowGPT: Knowledge Injection for Large Language Models
Qinggang Zhang1*, Junnan Dong1*, Hao Chen1, Daochen Zha2, Zailiang Yu3, Xiao Huang1 1The Hong Kong Polytechnic University, 2Rice University, 3Zhejiang Lab {qinggangg.zhang,hanson.dong}@connect.polyu.hk, sundaychenhao@gmail.com daochen.zha@rice.edu, yuzl@zhejianglab.com, xiaohuang@comp.polyu.edu.hk
@@ 1 Introduction
Large language models (LLMs) have surprised the world with their superior performance Kung et al. (2023); Zha et al. (2023), especially with the emergence of ChatGPT and GPT-4 OpenAI (2023). Nonetheless, LLMs are often criticized for their limited factual knowledge and propensity to produce hallucinations, wherein the model fabricates incorrect statements on tasks beyond their knowledge and perception Amaro et al. (2023); Shen et al. (2023); Gravel et al. (2023). Consider an ecological domain-specific question from OpenbookQA Mihaylov et al. (2018) in Figure 1. ChatGPT erroneously responds “energy” when asked about the portion of nutrients. This inaccuracy could stem from its potential lack of knowledge of carbs and their relationship to nutrients.
@@
Figure 1: A real-world question in OpenbookQA and a subgraph of ConceptNet. ChatGPT could effectively correct the answer given the scientific reasoning background in KG (blue: question concepts, red: candidate answers).
A promising avenue for addressing the above issue entails the integration of Knowledge Graphs (KGs) into LLMs. KGs Li et al. (2022a, b, 2023); Zhang et al. (2022a, 2023), such as Yago Suchanek et al. (2007), Freebase Bollacker et al. (2008), and ConceptNet Speer et al. (2017) represent relationships among real-world entities in a structured form as triples (head, relation, tail). The enormous factual knowledge stored in KGs holds the potential to significantly enhance the accuracy of LLMs’ responses. For instance, in Figure 1, ChatGPT could correct itself by leveraging the related background knowledge in ConceptNet Speer et al. (2017).
Recently, various white-box knowledge injection methods have been proposed to incorporate KGs into LLMs. These methods assume that we know everything about LLMs, such as model architectures and parameters. Many algorithms have been proposed to integrate KGs into LLMs Pan et al. (2023). Early studies directly concatenate the entities from KGs with textual sentences as the input to train LLMs based on cross-modal representation alignment Sun et al. (2020b); Liu et al. (2020). Later research incorporates KGs at an implicit level by either combining text representation and KG embedding with the attention mechanism Feng et al. (2020); Yasunaga et al. (2021); Dong et al. (2023a) or designing various fusion methods to integrate KGs and texts through a tailored encoder Zhang et al. (2019); Sun et al. (2020a).
@@
Unfortunately, many state-of-the-art LLMs are confined to be close-source in practical applications. For instance, ChatGPT and GPT-4 exclusively grant access through their APIs, which means we can only retrieve model responses by submitting textual inputs, with model specifics inaccessible. This lack of access prevents us from employing the aforementioned white-box knowledge injection techniques. Even though white-box approaches could be applied to open-source LLMs, such as BLOOM Scao et al. (2022) and LLaMA Touvron et al. (2023), they will often incur significant computation costs due to updating model weights Liu et al. (2022). Thus, we ask: Can we develop a knowledge injection framework that can efficiently and effectively integrate KGs into LLMs with APIs only?
@@
Achieving this goal is nontrivial because of two challenges in constructing model inputs (prompts). Identifying the most relevant knowledge is difficult. Real-world KGs often consist of millions of triples, whereas LLMs are typically restricted by limited input lengths (e.g., 2048 tokens for ChatGPT and 4096 tokens for GPT-4). Hence, careful selection of the most informative knowledge from KGs becomes essential. Effectively encoding KG knowledge is hard. It is observed that even minor variations in prompts conveying the same semantic meaning can yield drastically different responses from LLMs OpenAI (2023). As a result, a customized approach to encoding factual knowledge from extracted KGs for each question is often required to achieve the best performance.
@@
In this work, we propose KnowGPT, a knowledge injection framework for LLMs in question answering. To address challenge , we leverage deep reinforcement learning (RL) to extract paths from source entities mentioned in the question to the target entities within the potential answers. To encourage the agent to discover more informative paths, we devise a tailored reward scheme that promotes the reachability, context-relatedness, and conciseness of the extracted paths. Then, a policy network is trained to maximize the reward using training questions and applied to unseen questions. To tackle challenge , we introduce a prompt construction strategy based on Multi-Armed Bandit (MAB). Given several path extraction strategies and prompt templates, a MAB is learned to select the most effective combination for each question by balancing exploration and exploitation. The learned MAB is then applied to new questions to select path extraction strategies and prompt templates automatically. Our main contributions are:
@@
Figure 2: The overall architecture of KnowGPT. Given the question context and answer choices, we retrieve a question-specific subgraph from the real-world KG. Knowledge Extraction is first dedicated to searching for the most informative and concise reasoning background subject to the context. Then the prompt translation module is optimized to prioritize the combination of knowledge and formats subject to the given question context.
  • Formally define the problem of balck-box knowledge injection for LLMs, which integrates KGs into LLMs with model APIs only.
  • Propose KnowGPT, a general knowledge injection model to capture and translate knowledge from Knowledge Graphs (KGs) into prompts.
  • Instantiate KnowGPT upon two real-world KGs, ConceptNet Speer et al. (2017) and USMLE Yasunaga et al. (2021), with ChatGPT APIs. KnowGPT outperforms the state-of-the-art baselines on three QA benchmarks by a large margin. Notably, KnowGPT attains a 91.6% accuracy on the OpenbookQA leaderboard, which is comparable to human performance.
2 Related Work
2.1 Integration of KGs and LLMs
Leveraging KG to assist LLMs for NLP tasks has been intensively studied. Colake Sun et al. (2020b) presents a unified graph that combines a word graph with the given context and KG. QA-GNN Yasunaga et al. (2021) takes this further and embeds the question context as an entity in the joint graph. Recent studies utilize the attention mechanism to incorporate KGs into LMs, thereby enhancing comprehension and reasoning processes. MHGRN Feng et al. (2020) iteratively refines the question embeddings from pre-trained LLMs via graph neural networks during reasoning. However, they are limited to adapt to LLMs since most existing LLMs are either closed-sourced or over-expensive to be effectively tuned.
To alleviate this issue, recent research has shown various ways to prompt the KG to LLMs. One heuristic way to incorporate KGs and LLMs is to inject triples as the input Pan et al. (2023). ERNIE3.0 Sun et al. (2021) takes triples as token sequences and straightforwardly appends them to the given sentences. Another group of work dedicate to designing various ways to retrieve the knowledge. Mindmap Wen et al. (2023) transforms the relevant knowledge into mindmaps and prompt the LLMs. ChatRule Luo et al. (2023) extracts multiple relational paths from KGs and prompts the LLMs with natural language sentences. CoK Wang et al. (2023) activates the LLMs’ comprehension through chain-of-knowledge which prompts triples step-by-step. However, they lack the prompt format design which is laborious to be expanded to more LLMs and task scenarios. In this work, we develop a knowledge injection framework that values both knowledge retrieval and prompt construction. It automatically prioritize relevant knowledge and prompts the LLMs with the the most suitable prompt formats.
@@ 3 Problem Statement
We formally define the problem of black-box knowledge injection for LLMs in complex question answering. We represent each question as a question context $\mathcal{Q}=\{\mathcal{Q}_{s},\mathcal{Q}_{t}\}$, where $\mathcal{Q}_{s}=\{e_{1},e_{2},...,e_{m}\}$ is a set of $m$ source entities, and $\mathcal{Q}_{t}=\{e_{1},e_{2},...,e_{n}\}$ is a set of $n$ target entities. Following prior work Feng et al. (2020); Yasunaga et al. (2022), $\mathcal{Q}_{s}$ is extracted by concept recognition, and we assume it is given in our problem. Similarly, each target entity in $\mathcal{Q}_{t}$ is extracted from a corresponding candidate answer. We denote an LLM as $f$, a real-world KG as $\mathcal{G}$, which consists of triples (head entity, relation, tail entity), denoted as $(h,r,t)$. In our setting, we only have access to the APIs of $f$. However, we can employ open-source lightweight language models (not $f$), like Bert-Base Kenton and Toutanova (2019), to obtain text embeddings. Using the above notations, we describe our problem below. Given a question context $\mathcal{Q}$, an LLM $f$, and a KG $\mathcal{G}$, we aim to learn a prompting function $f_{\text{prompt}}(\mathcal{Q},\mathcal{G})$, which generates a prompt x that incorporates the context of $\mathcal{Q}$ and the factual knowledge in $\mathcal{G}$, such that the prediction of the LLM $f(\textbf{x})$ can output the correct answers for $\mathcal{Q}$.
@@ 4 KnowGPT Framework
Learning the prompting function $f_{\text{prompt}}(\mathcal{Q},\mathcal{G})$ involves two challenges, i.e., what knowledge should be used in $\mathcal{G}$, and how to translate the structured knowledge into prompts. To address these challenges, we present KnowGPT, which extracts subgraphs (paths) with deep RL and then constructs the prompt with MAB. An overview of our framework is shown in Figure 2.
4.1 Knowledge Extraction with Deep Reinforcement Learning
Intuitively, the relevant reasoning background lies in a question-specific subgraph $\mathcal{G}_{\text{sub}}$ that contains all the source entities $\mathcal{Q}_{s}$, target entities $\mathcal{Q}_{t}$, and their neighbors. An ideal subgraph $\mathcal{G}_{\text{sub}}$ is expected to have the following properties: $(i)$ $\mathcal{G}_{\text{sub}}$ encompasses as many source and target entities as possible, $(ii)$ entities and relations within $\mathcal{G}_{\text{sub}}$ exhibit a strong relevance to the question, and $(iii)$ $\mathcal{G}_{\text{sub}}$ is concise with little redundant information such that it can be fed into LLMs with limited lengths.
However, it is challenging to find such a $\mathcal{G}_{\text{sub}}$ since extracting a subgraph is NP-hard. To effectively and efficiently find a satisfactory $\mathcal{G}_{\text{sub}}$, we develop a tailored knowledge extraction method, named $\mathcal{P}_{\text{RL}}$, that employs deep RL to sample reasoning paths in a trial-and-error fashion. Specifically, we assume $\mathcal{G}_{\text{sub}}$ is constructed based on a set of reasoning paths $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{m}\}$, where each path $\mathcal{P}_{i}=\{(e_{i},r_{1},t_{1}),(t_{1},r_{2},t_{2}),...,(t_{|\mathcal{P}_{i}|-1},r_{|\mathcal{P}_{i}|},t_{|\mathcal{P}_{i}|})\}$ is a path in $\mathcal{G}$ starting from the $i$-$th$ source entity in $\mathcal{Q}_{s}$, and $|\mathcal{P}_{i}|$ is the path length. $\mathcal{G}_{\text{sub}}$ encompasses all the entities and relations appeared in $\mathcal{P}$. We model the sampling of each reasoning path as a Markov Decision Process (MDP) with state, action, transition, and reward, defined as follows.
@@ 4.1 Knowledge Extraction with Deep Reinforcement Learning
    where $\boldsymbol{e}_{t}$ and $\boldsymbol{e}_{target}$ are the embedding vectors of the current entity and the target entity. To get the initial node embeddings for entities extracted from the background KG, we adopt the approach proposed by the previous study Feng et al. (2020). Specifically, we transform knowledge from the KG into sentences and feed them into pre-trained LM to get node embeddings.
@@ 4.1.1 Context-relatedness Auxiliary Reward
where $\boldsymbol{c}$ is the embedding of context $\mathcal{Q}$ we obtained from a pre-trained LM Kenton and Toutanova (2019) and the embedding of path $\mathcal{P}_{i}$ is the average of the embeddings of all the entities and relations we have walked through till $i$, i.e., $Avg(\boldsymbol{e}_{source}+\boldsymbol{re}_{1}...+\boldsymbol{e}_{i})$, where $i\leq length(\mathcal{P}_{target})$. This step-by-step reward scheme provides rewards before the target is reached.
@@ 4.1.3 Training Policy Network
4.2 Structured Knowledge Prompts
@@
Suppose we have several path extraction strategies $\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{m}\}$ and several candidate formats =$\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},...,\mathcal{F}_{n}\}$. Each path extraction strategy $\mathcal{P}_{i}$ is a method for selecting a subgraph given a question context, such as the RL-based strategy discussed above. Every prompt template $\mathcal{F}_{j}$ represents a transformation of the triples within the subgraph into a natural language prompt.
@@
To capture context-aware correlation between questions and different combinations of knowledge and prompt formats, we formulate the selection mechanism of MAB with an expectation function $E(\cdot)$. It adaptively measures the potential expectation of a combination for different questions.
$E(\mathcal{Q}|\mathcal{P}\mathcal{F}_{(i)})=\boldsymbol{c}\times\bm{\alpha}_{(i)}+\beta_{(i)}.\vspace{1mm}$ (6)
@@
Empirically maximizing $\boldsymbol{c}\times\bm{\alpha}_{i}$ could encourage exploitation Chen et al. (2019); Dong et al. (2023b) for the best combination, we could effectively update $\bm{\alpha}_{(i)}$ via modeling the correlations between the context embedding of the anchor question ${\bf c}_{i}$ and all the previously selected contexts ${\bf C}_{(i)}$ for particular combination $\mathcal{P}\mathcal{F}_{(i)}$ in former $k$ steps, and the rewards $\textbf{r}_{pf}^{(i)}$ obtained from the selection of the current combination. Concretely, $\bm{\beta}^{(b)}$ is updated as: